OpenStudy (anonymous):

Assume that a & b can represent any real number: ^448√(a+b)^448=

jimthompson5910 (jim_thompson5910):

\[\large \sqrt[448]{(a+b)^{448}}=|a+b|\] This is using the fact that \[\large \sqrt[n]{x^n}=|x|\] where n is an even integer.